Shift report 20011017 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Oct 17 2001 - 23:56:00 EDT

  • Next message: Apache: "Shift report 20011018 00:00-08:00"

    ------------------------------------------------------------
    Date: Wed Oct 17 23:55:59 2001, Shift: 16:00-24:00
    Supervisor: KH, with: JC
    Log book#10 pages 25 - 25
    
    Shift Summary:
    ------------------------------------------------------------
    No beam during entire shift.
    There was a beam study during day shift.  It ended at
    16:00 with the intention to supply 110 bunch beam.
    They never succeeded.
    
    The only runs were pedestal runs.
    
    No detector problems noticed.
    
    Magnets are off (but enabled in expectation of beam)
    Anode voltages of TPCs and DCs are off.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Oct 17 2001 - 23:56:36 EDT